OpenStudy (anonymous):

For school, Georgia bought 7 notebooks and 3 folders. Let n represent the cost of each notebook and f represent the cost of each folder. Write an expression to determine the total cost for the school supplies. Then, calculate the cost if each notebook costs fifty cents and each folder costs two dollars. seven n plus three f; nine dollars and fifty cents seven n plus three f; fifteen dollars and fifty cents three n plus seven f; nine dollars and fifty cents three n plus seven f; fifteen dollars and fifty cents

OpenStudy (texaschic101):

7n + 3f = total cost 7(.50) + 3(2) = 3.50 + 6.00 = 9.50 Answer : A.